Prior to their phaseout in the 1980s, chemicals containing lead were commonly added to gasoline as anti‑knocking agents. A 8.773 g8.773 g sample of one such additive containing only lead, carbon, and hydrogen was burned in an oxygen‑rich environment. The products of the combustion were 9.549 g CO2(g)9.549 g CO2(g) and 4.886 g H2O(g)4.886 g H2O(g) .

Determine the empirical formula of the gasoline additive.

empirical formula:CHPb ?
Answer #1

mass of C in 9.549 g CO2 = (12 / 44) * 9.549 = 2.604 g

and

mass of H in 4.886 g H2O = (2 / 18) * 4.886 = 0.543 g

and

mass of Pb = 8.773 - (2.604 + 0.543) = 5.626 g

C : H : Pb

2.604 / 12 : 0.543 / 1 : 5.626 / 207.2

or

0.217 : 0.543 : 0.027

or

0.217 / 0.027 : 0.543 / 0.027 : 0.027 / 0.027

or

8 : 20 : 1

thus

empirical formula of the gasoline = C8H20Pb